OpenStudy (anonymous):

you are given a sample containing 7.5x10^6 cells/mL and told to prepare a sample of 3,000 cells/mL. What dilution factor is needed?

OpenStudy (anonymous):

I think it might work like this. (7.5/3) * 10^(6-3) = 2.5 * 10^3 = 2500

OpenStudy (anonymous):

well i know a dilution factor is the total volume/aliqout volume. i just don't know which would be the total volume or which the aliquot.

OpenStudy (anonymous):

I have never studied this concept, but I guess here, total volume may be equal to 7.5 * 10^6 , and aliquot volume may be equal to 3000.

OpenStudy (anonymous):

hmm...somebody said that in the math section too. imma go with it...this is an unfamiliar concept for me too. thanks =)

OpenStudy (anonymous):

Better to confirm I suppose.